# Latency-Aware Resource Allocation for Mobile Edge Generation and Computing via Deep Reinforcement Learning

### v1

@@ Abstract
Recently, the integration of mobile edge computing (MEC) and generative artificial intelligence (GAI) technology has given rise to a new area called mobile edge generation and computing (MEGC), which offers mobile users heterogeneous services such as task computing and content generation. In this letter, we investigate the joint communication, computation, and the AIGC resource allocation problem in an MEGC system. A latency minimization problem is first formulated to enhance the quality of service for mobile users. Due to the strong coupling of the optimization variables, we propose a new deep reinforcement learning-based algorithm to solve it efficiently. Numerical results demonstrate that the proposed algorithm can achieve lower latency than two baseline algorithms.
@@ I Introduction
In recent years, artificial intelligence generative content (AIGC) has gained widespread attention due to its powerful creative ability for a variety of content, such as images, videos, and music [1]. Several examples, including the generative pre-trained transformer (GPT) developed by OpenAI and the WaveNet developed by DeepMind, have shown great potential in enhancing communication performance for the next-generation wireless networks [2]. By deploying AIGC services at the network edge, lower latency and reduced communication overhead can be achieved for mobile users (MUs) [3]. Meanwhile, computing services remain crucial due to the increasing demand for handling computationally intensive tasks for MUs [4]. By processing task data at the network edge, the shortcomings of network congestion and long latency in conventional mobile computing systems can be addressed [5]. Hence, the integration of communication, computing, and generation services at the network edge is promising to address the heterogeneous requests in future wireless networks.
@@
Most previous works focus on the computing resource assignments or the AIGC service allocations in MEC and MEG, while the integration of MEC and MEG to provide heterogeneous services for MUs has not been exploited. To provide heterogeneous computation, AIGC, and vision enhancement (VE) services for MUs while improving the user experience, we aim to minimize the average latency of all MUs in a novel mobile edge generation computing (MEGC) system. A DRL-based latency-aware resource allocation algorithm is proposed to jointly optimize the bandwidth allocation, the backhaul transmit power, the computation resources, and the task offloading ratio. The superior latency performance of the proposed algorithm is verified by comparing to two benchmark algorithms.
@@ II System Model and Problem Formulation
To enable the heterogeneous task requirements, we consider an MEGC system that can provide both computation services and AIGC services, as illustrated in Fig. 1. The MEGC system consists of an edge server (ES) equipped with a powerful computing and generating unit, and three MUs with different task requests, including computation, AIGC, and VE, respectively. Let $U_{\mathsf{comp}}$, $U_{\mathsf{AIGC}}$, $U_{\mathsf{VE}}$ denote the MUs with computation task, AIGC task, and VE task, respectively. The serving time is index by $\mathcal{T}=\{1,\cdots,T\}$. Each MU requests one service during each time slot, and the frequency division multiple access (FDMA) technology is adopted in each time slot of request turn.The system model is shown in Fig 1.
@@ III DRL-based Resource Allocation Algorithm
Since the optimization variables of the bandwidth allocation $\bm{\alpha}$, the transmit power allocation of the ES $\bm{\beta}$, the computation resource of the ES $\bm{\omega}$, and the offloading ratio $\bm{\lambda}$ are highly coupled, the problem P1 is hard to solve by traditional optimization algorithms. Fortunately, DRL can effectively address complex resource allocation challenges by dynamically adapting decisions to changing environments that include varying channels and diverse task requests [12]. Therefore, we propose a DRL-based latency-aware resource allocation (LARA) algorithm for solving problem P1.
@@ IV Numerical Results
In this section, we test the performance of the proposed LARA algorithm by comparing it with two benchmarks: 1) fixed resource allocation (FRA) algorithm, where the optimization variables in P1 are fixed; 2) random resource allocation (RRA) algorithm, where the optimization variables in P1 are randomly selected in the feasible region. The distances between the ES and the MU $U_{\mathsf{comp}}$, the MU $U_{\mathsf{AIGC}}$, and the MU $U_{\mathsf{VE}}$, are $100,120,80$ meters, respectively. The transmit power of the MUs are $p_{\mathsf{comp}}=p_{\mathsf{VE}}=15$W, and the system bandwidth is $B_{\mathsf{off}}=B_{\mathsf{back}}=400$MHz. The transmit power of the ES is $P=15$W. The offloading task data volume of the MUs $U_{\mathsf{comp}}$ and $U_{\mathsf{VE}}$ follows a uniform distribution of $[1,8]$ Mbits. The noise power density is $N_{0}=-100$dBm/Hz. The GPT-4 is employed as the AIGC model [2], and the generation-related parameters $\xi$ and $\zeta$ are $9.97\times 10^{-14}$ and $5.73$, respectively. The generated data volume $D_{\mathsf{AIGC}}^{\mathsf{ES}}$ and $D_{\mathsf{VE}}^{\mathsf{ES}}$ are taken from $[2,16]$ Mbits according to specific AIGC and VE requests, respectively.
@@
Fig. 3 illustrates the training process of the proposed LARA algorithm across different learning rates (LRs). The orange curve shows that a smaller LR results in excessively slower updates in policy, increasing the possibility of falling into a sub-optimal policy. Hence, a smaller LR may not yield better outcomes. The blue curve demonstrates that an appropriate LR can facilitate the training process, achieving faster convergence performance. Fig. 4 shows the latency performance versus the training episodes for the three MUs. Despite each MU having distinct service requirements and varying communication and computation resources, the latency of each MU eventually converges. This observation validates the effectiveness and adaptability of the proposed LARA algorithm in tackling complex and heterogeneous resource allocation problems.
Figure 4: Latency versus the training episodes of different MUs: (a) the MU $U_{\mathsf{comp}}$, (b) the MU $U_{\mathsf{AIGC}}$, (c) the MU $U_{\mathsf{VE}}$.
Figure 5: The average latency vs. the optimization strategy.
Fig. 5 shows the latency of each MU and the total latency for all MUs under the proposed LARA algorithm and two benchmark algorithms. The proposed LARA algorithm can achieve the smallest latency among the three algorithms. The reason is that the proposed LARA algorithm is capable of adjusting the action according to the time-varying state information. Furthermore, each MU’s latency performance in the proposed LARA algorithm achieves the lowest latency compared to the corresponding MU in the benchmarks. The fairness of the proposed LARA algorithm to provide services to all MUs at the same time is thus verified. In addition, we can observe that the latency of the MU $U_{\mathsf{comp}}$ is the lowest among the three MUs. The reason is that the MU $U_{\mathsf{comp}}$ does not request AIGC or VE services at the ES, while the backhaul transmission latency for the computation results can be ignored. In contrast, the MUs $U_{\mathsf{AIGC}}$ and $U_{\mathsf{VE}}$ require the computation resources of the ES for the AIGC/VE services, and also require the backhaul transmissions for the results with an assignable amount of data volume.
@@ V Conclusion
This letter investigated the latency minimization problem in an MEGC system to provide computation, AIGC, and VE services for MUs. The bandwidth allocation, the backhaul transmit power, the computation resources, and the task offloading ratio, are jointly optimized. A novel DRL-based LARA algorithm is designed to solve the optimization problem. Finally, simulation results demonstrate that the proposed LARA algorithm outperforms the baseline algorithms in terms of the latency performance.

### v2

@@ Abstract
Recently, the integration of mobile edge computing (MEC) and generative artificial intelligence (GAI) technology has given rise to a new area called mobile edge generation and computing (MEGC), which offers mobile users heterogeneous services such as task computing and content generation. In this letter, we investigate the joint communication, computation, and the AIGC resource allocation problem in an MEGC system. A latency minimization problem is first formulated to enhance the quality of service for mobile users. Due to the strong coupling of the optimization variables, we propose a new deep reinforcement learning-based algorithm to solve it efficiently, Numerical results demonstrate that the proposed algorithm can achieve lower latency than several baseline algorithms.
@@ I Introduction
In recent years, artificial intelligence generative content (AIGC) has gained widespread attention due to its powerful creative ability for a variety of content, such as images, videos, and music [1]. Several examples, including the generative pre-trained transformer (GPT) developed by OpenAI and the WaveNet developed by DeepMind, have shown great potential in enhancing communication performance for the next-generation wireless networks [2]. By deploying AIGC services at the network edge, lower latency and reduced communication overhead can be achieved for mobile users (MUs) [3]. Meanwhile, computing services remain crucial due to the increasing demand for handling computationally intensive tasks for MUs [4]. By processing task data at the network edge, the shortcomings of network congestion and long latency in conventional mobile computing systems can be addressed [5, 6]. Hence, the integration of communication, computing, and generation services at the network edge is promising to address the heterogeneous requests in future wireless networks.
@@
Most previous works focus on the computing resource assignments or the AIGC service allocations in MEC and MEG, while the integration of MEC and MEG to provide heterogeneous services for MUs has not been exploited. To provide heterogeneous computation, AIGC, and vision enhancement (VE) services for MUs while improving the user experience, we aim to minimize the average latency of all MUs in a novel mobile edge generation computing (MEGC) system. A DRL-based latency-aware resource allocation algorithm is proposed to jointly optimize the bandwidth allocation, the backhaul transmit power, the computation resources, and the task offloading ratio. The superior latency performance of the proposed algorithm is verified by comparing to benchmark algorithms.
@@ II System Model and Problem Formulation
To enable the heterogeneous task requirements, we consider an MEGC system that can provide both computation services and AIGC services, as illustrated in Fig. 1. The MEGC system consists of an edge server (ES) equipped with a powerful computing and generating unit, and three kinds of MUs with different task requests, including computation, AIGC, and VE, respectively111 The MEGC system can be directly extended to the system with multiple MUs, where the MUs with identical tasks can be aggregated as a cluster. . et $U_{\mathsf{comp}}$, $U_{\mathsf{AIGC}}$, $U_{\mathsf{VE}}$ denote the MUs with computation task, AIGC task, and VE task, respectively. The serving time is index by $\mathcal{T}=\{1,\cdots,T\}$. Each MU requests one service during each time slot, and the frequency division multiple access (FDMA) technology is adopted in each time slot of request turn.The system model is shown in Fig 1.
@@ II-C Problem Formulation
Since the objective function of problem P1 is non-convex, problem P1 is a non-convex optimization problem. In addition, the optimization variables in P1 are highly coupled. Hence, problem P1 is hard to solve by traditional optimization algorithms. Fortunately, DRL can effectively address complex resource allocation challenges by dynamically adapting decisions to changing environments that include varying channels and diverse task requests [13]. Therefore, we propose a DRL-based latency-aware resource allocation (LARA) algorithm for solving problem P1.
@@ IV Numerical Results
In this section, we test the performance of the proposed LARA algorithm by comparing it with four benchmarks: 1) fixed resource allocation (FRA) algorithm, where the optimization variables in P1 are fixed; 2) random resource allocation (RRA) algorithm, where the optimization variables in P1 are randomly selected in the feasible region; 3) discrete action based resource allocation (DARA) algorithm, where the optimization variables in P1 are chosen from discrete actions, and the deep $\mathcal{Q}$-learning network is adopted; 4) exhaustive searching scheme (ESS), it discretizes the actions into multiple finely spaced discrete values and uses the exhaustive search algorithm to explore all possible solutions of these actions. By selecting the action that results in the minimum average time latency, the ESS effectively approaches the problem P1’s lower bound solution. Due to the adopted exhaustive search algorithm for selecting actions, the ESS has high complexity[15, 16]. The distances between the ES and the MU $U_{\mathsf{comp}}$, the MU $U_{\mathsf{AIGC}}$, and the MU $U_{\mathsf{VE}}$, are $100,120,80$ meters, respectively. The transmit power of the MUs are $p_{\mathsf{comp}}=p_{\mathsf{VE}}=15$W, and the system bandwidth is $B_{\mathsf{off}}=B_{\mathsf{back}}=400$MHz. The transmit power of the ES is $P=15$W. The offloading task data volume of the MUs $U_{\mathsf{comp}}$ and $U_{\mathsf{VE}}$ follows a uniform distribution of $[1,8]$ Mbits. The noise power density is $N_{0}=-100$dBm/Hz. The GPT-4 is employed as the AIGC model [2], and the generation-related parameters $\xi$ and $\zeta$ are $9.97\times 10^{-14}$ and $5.73$, respectively. The generated data volume $D_{\mathsf{AIGC}}^{\mathsf{ES}}$ and $D_{\mathsf{VE}}^{\mathsf{ES}}$ are taken from $[2,16]$ Mbits according to specific AIGC and VE requests, respectively.
@@
Figure 4: Latency versus the training episodes: (1) the MU $U_{\mathsf{comp}}$, (2) the MU $U_{\mathsf{AIGC}}$, (3) the MU $U_{\mathsf{VE}}$.
Fig. 4 illustrates the training process of the proposed LARA algorithm under different learning rates (LRs) for both large network (LN) and small network (SN) structures. The LN-based LARA comprises $5$ hidden layers, and the maximum number of neurons in each layer is $4,096$. In contrast, the SN-based LARA consists of $4$ hidden layers, and each layer has a maximum number of $600$ neurons. It is evident that the LN-based LARA achieves better convergence rewards and maintains stability during training. The red curve indicates that when the LARA has large network structure, a smaller LR can more closely approximate the optimal policy. Conversely, the orange curve demonstrates that when the network has inadequate hidden layers and neurons, a smaller LR may result in slow policy updates, leading to a suboptimal strategy. Fig. 4 shows the latency performance versus the training episodes for the three MUs. Despite each MU having distinct service requirements and varying communication and computation resources, the latency of each MU eventually converges. This observation validates the effectiveness and adaptability of the proposed LARA algorithm in tackling complex and heterogeneous resource allocation problems.
Figure 5: The average latency vs. different algorithms.
Figure 6: The total latency vs. different algorithms.
Fig. 6 and Fig. 6 show the latency of each MU and the total latency for all MUs under the proposed LARA algorithm and two benchmark algorithms. The proposed LARA algorithm can achieve the smallest latency among all the algorithms, and has near performance to the ESS algorithm. The reason is that the proposed LARA algorithm is capable of adjusting the action according to the time-varying state information in a continuous action selection manner. Furthermore, each MU’s latency performance in the proposed LARA algorithm achieves the lowest latency compared to the corresponding MU in the four benchmarks, and also close to the ESS. The fairness of the proposed LARA algorithm to provide services to all MUs at the same time is thus verified. In addition, we can observe that the latency of the MU $U_{\mathsf{comp}}$ is the lowest among the three MUs. The reason is that the MU $U_{\mathsf{comp}}$ does not request AIGC or VE services at the ES, while the backhaul transmission latency for the computation results can be ignored. In contrast, the MUs $U_{\mathsf{AIGC}}$ and $U_{\mathsf{VE}}$ require the computation resources of the ES for the AIGC/VE services, and also require the backhaul transmissions for the results with an assignable amount of data volume.
@@ V Conclusion
This letter investigate the latency minimization problem in an MEGC system to provide computation, AIGC, and VE services for MUs. The bandwidth allocation, the backhaul transmit power, the computation resources, and the task offloading ratio, are jointly optimized. A novel DRL-based LARA algorithm is designed to solve the optimization problem. Finally, simulation results demonstrate that the proposed LARA algorithm outperforms the baseline algorithms in terms of the latency performance. Besides, the selections of the LR and network structure are also analyzed.